Dear colleagues,
Here is a summary of today's activities at CSBF.
- Jim re-clamped the flat cable connector on the MXR box. The error message does not show up after that in test mode whole day. But as soon as the muon run in flight mode started in a few minutes, it came back. We will see if this disappears as time goes on as in last two nights.
- Thermal/Vacuum chamber was shown by Joseph. The specification is summarized and uploaded on the black board.
- ZLo puzzle solved by replacing the LVDS cable, which was used from at least last flight. Now the ZLo trigger behaves as we expected.
- Jim told me that the ZLo trigger generated in the TCD trigger circuit will be sent to the MT independent from the MT confirmation as we expected from the ZLo trigger behavior. But the digitization starts only when the MT sends a confirmation (decided based on the trigger logic) trigger within 1 or 1.5 micro seconds.
- Debugging of the reduced hsk data being worked by Prashant.
- Data copying for the last 3 day's CPT data are done.
- SangEun has started working on the SIP interface coding.
- Light shielding of box +X, -Y boxes are done. So all boxes are now light tight.
- Cabling in -X sides, where Switch box, SFC, Command, CAL Trigger are located, are cleaned up. So only -X side needs cleaning up.
- Dave and Fritz came around 5:30PM. They will start work from tomorrow.
- Power cable and 2 SSDs are sent from UMD. (Larry/Dr. Kim)
Here are things to do.
- TCD mxr debugging.
- S3 tuning.
- CD housekeeping conversion factor in the housekeeping table.
- CD integration.
- Repair the MUX cable connector (9-pin D-sub connector) for the Chercam HSK.
- CDAQ remaining works including reduced housekeeping. (Prashant)
- Clean up -X side cables.
- Test of temperature sensors for CDM.
- Power cables for CDM shipping on Monday. (Larry) etc...
Please let us know if you have any questions/comments/suggestions.
Cheers, MooHyun.
MooHyun, July 14, 2008



